You are evaluating the following mutually exclusive projects for your firm, whose cost of capital is 14%, and all dollar amounts are in millions.

1. Verify the NPV and IRR of each project.

2. What is your recommendation?

Project

Required Return

Life

IO

NCF1-n

NPV

IRR

Alpha

12%

10 years

$50

$20

$63

38.45%

Beta

8%

5

$50

$25

$49.82

41.04%

EAABeta = $12.48 million > EAAAlpha = $11.15 million so

Accept B and Reject A Is the answer- Please explain why
Answer #1

Answer 1)

Project Rate Life IO NCF NPV IRR
Alpha 12% 10 $50 $20 $63.00 PV(12%,10,-20)-50 38%
Beta 8% 5 $50 $25 $49.82 PV(8%,5,-25)-50 41.04%

Answer 2)

Here, recommendation should be based out on the Equivalent Annual Annuity Approach.

EAA = (r x NPV) / (1 - (1 + r)-n )

for Project Alpha EAA = (0.12*63)/(1-(1.12^-10)) = $ 11.150

for Project Beta EAA = (0.08*48.82)/(1-(1.08^-5)) = $ 12.48

As EAA of Beta > EAA of Alpha ,

So, Accept B and Reject A